Determine the stability or instability of all solutions of the following systems of differential equations

.= 1 -2 0 2 101 0 1 0 0 0 2“ OOON 10 0 -2 0 1
Answer #1

Solution 1-2 0 0 0 0 0 0 0 -2 1 2 0 x = - where AX To 210 -2 0 0 1 A = Loo -2 het eigenvalue of A is 1 A-1) = 0 - 13 | -2 9 oi hope it will help you...

For any quary comment please...